a charged particle of mass 2 mg remains freely in air in an elecrric field of strength 4 N/C directed upward. calculate the charge and determine its nature (g=10m/s2)

Grade:11

3 Answers

Raman Mishra
7 years ago
The given electric field should exert an upward force on the particle so that it cancels the gravitational force which acts on the particle in the downward direction.
Since the electric field is directed upwards, therefore the particle should be positively charged so that an upward force acts on it due to the electric field.
Now, net force on the particle = Fnet = 0
Taking the upward direction as positive, we have:
                   Fnet = Fe – Fg = 0
               => qE – mg = 0
               => q = mg/E
   Substituting m=2\times 10^{-3}kg 
                      E = 4N/C and g = 10m/s2 , we get:
       q=5\times 10^{-3}C   Ans.
Xigroyr6
6 years ago
The given electric field should exert an upward force on the particle so that it cancels the gravitational force which acts on the particle in the downward direction.Since the electric field is directed upwards, therefore the particle should be positively charged so that an upward force acts on it due to the electric field.Now, net force on the particle = Fnet = 0Taking the upward direction as positive, we have: Fnet = Fe – Fg = 0 => qE – mg = 0 => q = mg/E Substituting m=2\times 10^{-3}kg E = 4N/C and g = 10m/s2 , we get: q=5\times 10^{-3}C
Krish Gupta
3 years ago
Pretty easy, just do force balance and you are done.
The given electric field should exert an upward force on the particle so that it cancels the gravitational force which acts on the particle in the downward direction.
Since the electric field is directed upwards, therefore the particle should be positively charged so that an upward force acts on it due to the electric field.
Now, net force on the particle = Fnet = 0
Taking the upward direction as positive, we have:
                   Fnet = Fe – Fg = 0
               => qE – mg = 0
               => q = mg/E
   Substituting m=2\times 10^{-3}kg 
                      E = 4N/C and g = 10m/s2 , we get:
       q=5\times 10^{-3}C   Ans.